OpenStudy (anonymous):

The circle formed by the blades of a fan has a circumference of fifty-six and four-sevenths inches. What is its diameter? Use twenty-two sevenths for pi. twelve inches eighteen inches twenty-one inches twenty-six inches PLEASE HELP I WILL FAN AND MEDAL

OpenStudy (anonymous):

Please help me #ineedhelp

OpenStudy (anonymous):

@ineedhelp

OpenStudy (anonymous):

@shiraz14 can you help???

OpenStudy (shiraz14):

Circumference of a circle, C = πd Given C = 56 4/7 & using π=22/7, we have: 396/7 = 22d/7 d = 18 in

OpenStudy (anonymous):

Oh Ok Man You Da Best Bro Thank You

OpenStudy (shiraz14):

You're welcome